Razor:

Simplify 1/8^2 Express your answer both with a positive exponent and in standard form

Razor:

@JustSaiyan @Elsa213

Razor:

@Eiwoh2

Eiwoh2:

Simplified, taht's 1/64.

Eiwoh2:

Wait no...

JustSaiyan:

Even though this isn't particularly difficult, I stand by my words of only doing my own Mathematics.

Eiwoh2:

Calculated it wrong.... x'D

Razor:

.... XD

Eiwoh2:

No, yeah, calculated, that'll be 1/64. As a decimal, 0.015625. ^^

Razor:

That's in standard form correct? and with a positive....exponent XD

Eiwoh2:

It's a fraction, but yes. :,)

Razor:

mkay thanks Howie... you just helped Howie jr/Razor finish an assignment XD

Eiwoh2:

Youch. That's what we're here for. x'D

dude:

\(\frac{1}{8^2}\)? Do \(8^2\) and leave it as a fraction to keep the positive exponent

Razor:

DOUBLE THANKS XD

Razor:

wait there was a mistake in my question XD it should be negative 2 exponent..

Eiwoh2:

Y O U C H

Razor:

so 1/8^-2

Eiwoh2:

Well, that seems more easier, i guess. Same way that Elsa put up there, and it would equal to 64, i believe.

Razor:

dude*

Razor:

64 negative or just positive?

Eiwoh2:

Just positive.

dude:

Oh, in that case \(\frac{1}{8^{-2}}\) becomes \(8^2\)

dude:

Yeah

Eiwoh2:

8^2 = 64

Razor:

Got it ... thanks lol

Eiwoh2:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Eiwoh2 Youch. That's what we're here for. x'D \(\color{#0cbb34}{\text{End of Quote}}\)